Estimate the increase in pressure in(Psi) required to decrease a unit volume of Mercury by 0.1%

1
Expert's answer
2022-01-31T17:07:02-0500

Increase in pressure in(Psi)


"E_v=-\frac{dp}{\frac{dV}{V}}"


"dp= -\frac{E_v\times dV}{V} =-\frac{4.14\times 10^6}{-0.001}=4.14\times10^3 psi"